a) Find a positive angle less than 2𝜋 that is coterminal with 16𝜋 3 radians.



b) Find the reference angle for 16𝜋 3 radians.


1
Expert's answer
2021-03-30T07:50:02-0400

Subtract "2\pi" from "16\pi"


"\Rightarrow" "\dfrac{16\pi}{3}−
2
π"

The resulting angle of"\dfrac{10\pi}{3}"  is positive, less than "2\pi" and coterminal with "\dfrac{16\pi}{3}"

"\dfrac{10\pi}{3}"

Subtract"2\pi" from "\dfrac{10\pi}{3}"

"\dfrac{10\pi}{3}-2\pi"

The resulting angle of "\dfrac{4\pi}{3}" is positive, less than and coterminal with "\dfrac{16\pi}{3}"

"\dfrac{4\pi}{3}"

Since the angle "\pi" is in the third quadrant subtract "\pi" from "\dfrac{4\pi}{3}"

"\Rightarrow \dfrac{\pi}{3}"